### v1

# Towards Distillation-Resistant Large Language Models: An Information-Theoretic Perspective

Abstract
Proprietary large language models (LLMs) embody substantial economic value and are generally exposed only as black-box APIs, yet adversaries can still exploit their outputs to extract knowledge via distillation. Existing defenses focus exclusively on text-based distillation, leaving the important logit-based distillation largely unexplored. In this work, we analyze this problem and present an effective solution from an information-theoretic perspective. We characterize distillation-relevant information in teacher outputs using the conditional mutual information (CMI) between teacher logits and input queries conditioned on ground-truth labels. This quantity captures contextual information beneficial for model extraction, motivating us to defend distillation via CMI minimization. Guided by our theoretical analysis, we propose learning a transformation matrix that purifies the original outputs to enhance distillation resistance. We further derive a CMI-inspired anti-distillation objective to optimize this transformation, which effectively removes distillation-relevant information while preserving output utility. Extensive experiments across multiple LLMs and strong distillation algorithms demonstrate that the proposed method significantly degrades distillation performance while preserving task accuracy, effectively protecting models’ intellectual property.
Machine Learning, ICML
1 Introduction
Large language models (LLMs) have recently demonstrated remarkable capabilities and have been deployed in a variety of practical applications, such as multi-agent systems (Wu et al., 2024; Du et al., 2024). Training such models demands substantial resources, including large-scale datasets, extensive computational budgets, and significant human effort, rendering them highly valuable intellectual property.
Figure 1: An illustration of the designed defense algorithm. We propose to learn a transformation matrix that post-processes output logits. By minimizing the conditional mutual information $\mathcal{I}(X;Z|Y)$, we successfully reduce the information beneficial for knowledge distillation within the teacher logits.
To safeguard their interests, commercial service providers typically isolate their proprietary models and only expose them through APIs. However, malicious attackers can exploit the model output texts or logits to perform knowledge distillation (KD) (Hinton et al., 2015; Kim and Rush, 2016; Gu et al., 2024), thus cheaply extracting model capabilities without any access to the weights or architecture.
To defend against such theft, early work has proposed LLM watermarking (Kirchenbauer et al., 2023; Ye et al., 2025; Liang et al., 2026) as a mitigation strategy, yet they offer only post-hoc verification and cannot directly prevent the extraction of proprietary knowledge. More recent efforts attempt to perturb the token distribution and modify model outputs to hinder KD (Savani et al., 2025; Li et al., 2025). However, these methods focus exclusively on text-based distillation, leaving the more widely discussed logit-based distillation largely unaddressed. Notably, a model’s output logits reveal the full probability distribution over tokens, conveying substantially richer information than the sampled token labels. Consequently, defending against such distillation is particularly challenging. This difficulty is further exacerbated by a growing body of work (Wang et al., 2021; Gu et al., 2024; Wang et al., 2025) that strengthens logit-based KD by more fully exploiting the rich information in these distributions. To address this challenge, we introduce an information-theoretic measure to quantify the distillation-relevant information in teacher logits and propose a transformation strategy to mitigate knowledge leakage.
Fundamentally, LLMs perform a classification task to predict the next token for an ongoing sequence. The information in an output vector can be interpreted in two aspects: (1) the predicted label for the current input (i.e., the token corresponding to the maximum logit), and (2) additional contextual information beyond the label (Ye et al., 2024). Although different prefixes can yield the same next predicted token, their logits differ substantially due to varying contexts in semantics and syntax. Particularly, logit-based KD exploits such contextual information to transfer richer knowledge to the student model (Gu et al., 2024).
To quantify the supervisory guidance from teacher outputs during distillation, we introduce conditional mutual information (CMI) as a principled measure of the conveyed information. Formally, let $X,Y,Z$ denote the input sequence, the ground-truth label, and the teacher logits for the given $X$, respectively. Then $\mathcal{I}(X;Z|Y)$ measures how much information $Z$ provides about $X$ conditioned on $Y$. This precisely aligns with the paradigm of logit-based KD that exploits information encoded in teacher logits beyond the label to enhance student performance. Therefore, a natural defense is to purify the teacher logits to minimize the conditional mutual dependence between $X$ and $Z$. To achieve this goal, we propose learning a transition matrix $\mathcal{M}$ that post-processes the output logits (see Figure 1), which we theoretically validate to reduce $\mathcal{I}(X;Z|Y)$. To effectively learn this matrix and decrease CMI, we further expand $\mathcal{I}(X;Z|Y)$ and decompose the minimization into two sub-problems based on the Information Bottleneck (IB) principle (IB开盒): (1) increasing the dependency between the logits $Z$ and the ground-truth label $Y$, and (2) weakening the correlation between the input $X$ and the logits $Z$. Guided by this analysis, we devise an optimization objective with two corresponding loss terms. We first derive a theoretical upper bound for the first subproblem, which formally justifies that minimizing the cross-entropy between $Z$ and $Y$ is a practical surrogate objective. The latter sub-task aims to reduce the relevance of $X$ and $Z$, yet directly computing or estimating this term for LLMs is intractable. To circumvent this difficulty, we note that under the KD scenario, the information in $Z$ about $X$ is ultimately leveraged through gradients propagated to the student during optimization. In this sense, these gradients serve as the practical carrier of the guidance provided by teacher logits. Building on this insight, we employ a surrogate student model to simulate distillation and maximize the angular deviation between the gradients induced by the transformed logits $Z^{\prime}$ and those from the original $Z$. By iteratively optimizing the transformation $\mathcal{M}$ with the proposed loss function, we effectively reduce the distillation-beneficial contextual information in teacher outputs, thereby strengthening their resistance against knowledge extraction attacks.
In summary, our contributions are as follows:
  • To the best of our knowledge, we are the first to defend LLMs against logit-based KD. We analyze the problem from an information-theoretic perspective and propose learning a matrix that post-processes output logits to minimize their CMI with the input.
  • Guided by theoretical analysis, we design two complementary loss functions to optimize the transformation matrix, which effectively removes the contextual information within the logits that could be exploited by attackers while preserving output utility.
  • Extensive experiments across various LLMs on three widely used datasets against four strong distillation algorithms demonstrate the outstanding defense performance of the proposed method.
2 Related Work
2.1 Knowledge Distillation for Large Language Models
Knowledge Distillation (KD) is a widely adopted paradigm for transferring knowledge from high-capacity teacher models to smaller and less capable student models. For large language models (LLMs), KD is usually realized through two paradigms: sequence-level (text-based) distillation and logit-based (response-level) distillation.
Sequence-level knowledge distillation trains the student model using texts generated by the teacher as hard-label supervision, treating the teacher-produced outputs as pseudo-labels and optimizing the student via standard maximum-likelihood estimation (Kim and Rush, 2016). In contrast, besides the SFT loss based on ground-truth labels, logit-based KD additionally uses the teacher’s soft output distributions to distill fine-grained predictive information by aligning the teacher and student distributions. The pioneering logit-based KD method (Hinton et al., 2015) uses the Kullback-Leibler divergence (KLD) to measure the discrepancy between the student and the teacher distributions. Reverse KLD (RKLD) employs a reverse version of KL divergence by changing the reference distribution in KL from the teacher to the student, allowing the student to concentrate its limited capacity on the teacher’s high-probability regions (Gu et al., 2024). AlphaNet (Wang et al., 2021) introduces $\alpha$-divergence, providing a tunable measure between KLD and RKLD to penalize over-estimation and under-estimation in KD. ABKD (Wang et al., 2025) further proposes a novel $\alpha\text{-}\beta$-divergence to balance hardness-concentration and confidence-concentration, generalizing FKLD and RKLD to a broader family of divergences.
2.2 Protecting LLMs from Knowledge Distillation
The great success of KD poses a significant threat to the protection of intellectual property in LLMs. To mitigate this issue, watermarking (Kirchenbauer et al., 2023; Liang et al., 2026) seeks to embed an identifiable and verifiable signature into a model’s behavior or parameters, enabling the model owner to verify their ownership. Meanwhile, Fingerprinting (Ye et al., 2025; Xu et al., 2024) assigns a unique and traceable identity to each distributed model instance, enabling the model owner to attribute leaked or misused models to their source. Despite their usefulness, both approaches are inherently post hoc, indicating that they facilitate the identification of model stealing after it has occurred, but cannot prevent the extraction process itself.
Another research line attempts to directly solve this threat by modifying the token distribution during decoding to obtain generated texts that are resilient to distillation. The pioneering work (Savani et al., 2025) proposes an anti-distillation decoding strategy that actively poisons reasoning traces by injecting a learned perturbation into the token-level probability distribution. Concretely, a proxy student first computes gradients on a small hold-out set of teacher-generated sequences, yielding a fixed poisoning direction $\nabla\ell$. During inference, the teacher performs two forward passes for each candidate token with proxy-student parameters perturbed by $\pm\eta\nabla\ell$, and estimates the resulting finite-difference logit shift. This shift is then integrated into the teacher’s logits to form an adjusted sampling distribution. Sampling from this modified distribution implicitly steers any downstream student trained on these traces toward higher loss, thereby degrading the effectiveness of distillation. DOGe (Li et al., 2025) adopts a more direct approach that maximizes the divergence between teacher and student output distributions to make the teacher’s predictions harder to imitate. Specifically, DOGe fine-tunes the teacher’s final linear layer via the negative KL-divergence between the temperature-scaled softmax of the teacher logits and the detached softmax of a proxy student, enlarging the teacher–student discrepancy.
However, existing methods limit their scope in the sequence-level distillation scenarios, leaving the defense against the more widely discussed logit-based KD algorithms largely unexplored. To address this gap, this work makes an early step in investigating this issue from an information-theoretic perspective and proposes learning a transformation matrix as an effective post-processing solution.
3 Preliminary
3.1 Notation
Let $\mathcal{V}$ be the vocabulary and $|\mathcal{V}|<\infty$. Let $X\in\mathcal{V}^{*}$ be the input sequence consisting of the user’s query and the ongoing generated response, and $Y\in\mathcal{V}$ be the ground-truth label, i.e., the next token for the current input $X$.
Let $f:\mathcal{V}^{*}\to\mathbb{R}^{|\mathcal{V}|}$ be the large language model, and $Z=f(X)$ be the logit vector for current input $X$. The next token prediction is sampled from the distribution $\sigma(Z)$, where $\sigma(\cdot)$ denotes the softmax function.
3.2 Knowledge Distillation
Given a dataset $\mathcal{D}$ consisting of prompt-answer pairs $(q,o)$, KD algorithms repeatedly query a teacher model $f_{\text{T}}(\cdot)$ with $(q,o)$ sampled from $\mathcal{D}$ and train a student model $f_{\text{S}}(\cdot)$ by mimicking the teacher’s output. Formally, the knowledge distillation loss can be expressed as:
$\mathcal{L}_{\text{KD}}\triangleq\frac{1}{l}\sum_{t=1}^{l}\text{Dist}\Big(\sigma\big(f_{\text{T}}(q,o_{<t})\big)\Big\|\sigma\big(f_{\text{S}}(q,o_{<t})\big)\Big),$ (1)
where $l$ denotes the length of the ground-truth answer $o$, and $\text{Dist}(\cdot\|\cdot)$ denotes the distance metric such as Kullback–Leibler divergence (KLD). In general, the objective of logit-based distillation algorithms can be formulated as:
$\mathcal{L}_{\text{total}}=(1-\alpha)\cdot\mathcal{L}_{\text{SFT}}+\alpha\cdot\mathcal{L}_{\text{KD}},$ (2)
where $\alpha$ is a factor set to $0.5$ by default, and $\mathcal{L}_{\mathrm{SFT}}$ denotes the supervised fine-tuning (SFT) loss, i.e., the negative log-likelihood of the predictive probability of the ground-truth label. Since the logits convey more contextual information than the predictive label alone, KD typically outperforms standard SFT on many downstream tasks. Based on this formulation, various improved algorithms have been proposed, including novel designs of distance metrics such as reverse KLD (Gu et al., 2024) and AB-KLD (Wang et al., 2025).
3.3 Information Bottleneck
The Information Bottleneck (IB) theory (IB提出; IB深度; IB开盒) highlights that a well-trained model should compress redundant information from inputs while preserving key information for tasks. The IB objective is formulated as:
$\mathcal{L}_{\text{IB}}=\mathcal{I}(X;Z)-\beta\cdot\mathcal{I}(Z;Y),$ (3)
where $\mathcal{I}(\cdot;\cdot)$ is mutual information, and $\beta>0$ is a hyperparameter. A lower $\mathcal{L}_{\text{IB}}$ indicates a better performance and higher robustness of a model. Note that $Z$ in IB represents the intermediate representation during inference. In this work, we instantiate $Z$ as the logit vector.
Figure 2: An illustration of the proposed defense framework. As shown in the left subfigure, we introduce a surrogate student model as an adversary and simulate the KD process. We compute the KD loss $\mathcal{L}_{\text{total}}$ and extract its gradients $\bm{g}$ with respect to the student parameters to quantify the guidance signal carried by the teacher outputs. The right subfigure explains the loss computation used for training the transformation matrix. To diminish context information in teacher logits, we design to minimize the KLD between the transformed logits $Z^{\prime}$ and the ground-truth label $Y$, while simultaneously maximizing the angular deviation between the gradients $\bm{g}$ and $\bm{g}^{\prime}$ calculated by the original and transformed logits, respectively.
4 Method
In this section, we first illustrate the rationality and benefits of minimizing CMI as a defense objective and present a theoretically-grounded post-processing strategy via a logit transformation matrix. Guided by theoretical insights, we design an optimization objective consisting of two complementary loss terms to train the matrix. An overview of the proposed method is illustrated in Figure 2.
4.1 Conditional Mutual Information-Based Defense
The CMI $\mathcal{I}(X;Z|Y)$ measures how much information $Z$ reveals about $X$ given the ground-truth label $Y$. The additional contextual information that $Z$ carries beyond the label aligns exactly with the goal of logit-based distillation, as shown in Equation 1. Therefore, we propose to defend teacher knowledge by minimizing $\mathcal{I}(X;Z|Y)$, preventing such contextual information from being extracted. This objective can also be interpreted through entropy and uncertainty. Specifically, $\mathcal{I}(X;Z|Y)=\mathcal{H}(X|Y)-\mathcal{H}(X|Z,Y)$, which quantifies the reduction in uncertainty of $X$ when $Z$ is additionally observed. For a fixed training dataset, $\mathcal{H}(X|Y)$ remains constant. Hence, minimizing $\mathcal{I}(X;Z|Y)$ is equivalent to maximizing $\mathcal{H}(X|Z,Y)$, i.e., making $Z$ contain as little information about $X$ as possible to increase the uncertainty about $X$ given the label $Y$.
Remark 1.
As an intuitive extreme case, if $Z$ ideally degenerates to the one-hot label $Y$, i.e., it contains no contextual information beyond the category, then logit-based knowledge distillation collapses to standard label-only supervised fine-tuning. In this case, the CMI attains its minimum value $\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Y|Y)=0$.
To reduce the contextual information in teacher logits, we propose a lightweight post-processing strategy that further modifies output logits while keeping the teacher completely frozen. Concretely, we introduce a learnable matrix $\mathcal{M}$ to transform the original teacher output into a refined one. Before delving into the full training pipeline, we show that such a transformation admits an intriguing information-theoretic property that formally guarantees that the refined logits become more context-compressed:
Theorem 2.
Let $X\in\mathcal{V}^{*}$ be the current input, $Y\in\mathcal{V}$ be the ground-truth next token, $Z=f(X)\in\mathbb{R}^{|\mathcal{V}|}$ be the teacher logits, $\mathcal{M}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ be the logit transformation matrix, and $Z^{\prime}=\mathcal{M}\cdot Z$ be the refined logits. Then
$Y\to X\to Z\to Z^{\prime}$
forms a Markov chain, and their CMI values satisfy
$\mathcal{I}(X;Z|Y)\geq\mathcal{I}(X;Z^{\prime}|Y).$
Detailed proofs can be found in Appendix A. Theorem 2 ensures that matrix transformations are universally beneficial for reducing CMI, which facilitates the compression of contextual information exploited by attackers. To further reduce CMI while preserving the utility of outputs, we additionally design an effective optimization objective to better guide the training of the transformation matrix.
4.2 Transformation Matrix Optimization
Next, we analyze the proposed matrix training objective through the lens of the Information Bottleneck (IB) principle to resolve the optimization problem. In particular, the following property reveals that minimizing CMI actually induces an IB objective.
Theorem 3.
The conditional mutual information $\mathcal{I}(X;Z|Y)$ is a special case of Information Bottleneck objective with $\beta=1$ in Equation 3, i.e.,
$\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Z)-\mathcal{I}(Z;Y).$ (4)
A detailed derivation is provided in Appendix A. This property is particularly appealing as it reveals that minimizing $\mathcal{I}(X;Z|Y)$ inherits the benefits of the Information Bottleneck, balancing compression of redundant information in $X$ while preserving task-relevant information in $Y$. The above characterization naturally decomposes the objective into two complementary sub-tasks:
  • Minimizing $\mathcal{I}(X;Z)$. This compresses redundant contextual information in the teacher logits $Z$ that originates from the input $X$, weakening the dependency between inputs and outputs. As a consequence, $Z$ becomes less informative about $X$, which enhances the resistance against KD algorithms.
  • Maximizing $\mathcal{I}(Z;Y)$. This preserves task-relevant information and maintains the model’s utility. Notably, it encourages $Z$ to concentrate toward its class centroid (i.e., increasing the correct label’s logit while suppressing others), which also eliminates contextual information and brings defense benefits.
Training Loss Design. Motivated by the two sub-tasks, we design two complementary loss functions to achieve the CMI minimization. We begin with analyzing the second term $\mathcal{I}(Z;Y)$, which can be expanded into
$\mathcal{I}(Z;Y)=\mathcal{H}(Y)-\mathcal{H}(Y|Z).$ (5)
Note that $\mathcal{H}(Y)$ is constant, thus maximizing $\mathcal{I}(Z;Y)$ is equivalent to minimizing the conditional entropy $\mathcal{H}(Y|Z)$, which measures the residual uncertainty about the ground-truth label $Y$ after observing logits $Z$. Although directly computing $\mathcal{H}(Y|Z)$ is intractable, we can derive a consistent upper bound as follows.
Lemma 4.
[…]
$\mathcal{H}(p,\hat{p})=\mathcal{H}(Y|Z)+\mathbb{E}_{Z}\left[\mathbb{KL}\Big(p(\cdot|Z)\Big\|\hat{p}(\cdot|Z)\Big)\right],$ (6)
which further implies the upper bound
$\mathcal{H}(p,\hat{p})\geq\mathcal{H}(Y|Z).$ (7)
Remark 5.
[…]
For maximizing $\mathcal{I}(X;Z)$, previous studies on image classification typically compute its value (Ye et al., 2024) or estimate it via variational upper bounds (IB变分). However, directly estimating $\mathcal{I}(X;Z)$ for autoregressive LLMs is fundamentally infeasible, since both $X$ and $Z$ are in high-dimensional spaces, and their underlying distributions are difficult to approximate. To handle this challenge, we note that in KD scenarios, the information of $X$ residing in $Z$ is ultimately exploited through how $Z$ instructs the student’s output, which in practice materializes in the form of gradients that update the student’s parameters during training. Hence, the learning signals conveyed by the gradients serve as the effective carrier of the information of $X$ residing in $Z$, which motivates the following proposition:
Proposition 6.
Let $\theta$ denote the parameters of a surrogate student model $f_{\textnormal{S}}(\cdot)$. Under the standard KD setting, the information in $Z$ about $X$ that can be exploited during distillation is operationally conveyed through the gradients of the loss with respect to $\theta$, i.e.,
$\bm{g}$ $=\nabla_{\theta}\frac{1}{l}\sum_{t=1}^{l}\Big[-(1-\alpha)\cdot\log\sigma\big(f_{\textnormal{S}}(q,o_{<t})\big)_{o_{t}}$ (8)
$+\alpha\cdot\mathbb{KL}\Big(\sigma\big(f_{\textnormal{T}}(q,o_{<t})\big)\Big\|\sigma\big(f_{\textnormal{S}}(q,o_{<t})\big)\Big)\Big].$
Based on the gradients $\bm{g}$, we can characterize the information within teacher logits $Z$ that is transmitted to the student during distillation. The optimization objective is accordingly reformulated as obtaining a transformed distribution $Z^{\prime}$ whose induced gradients $\bm{g}^{\prime}$ contain as little valid information as possible. Since the original $\bm{g}$ represents the correct and informative update direction under KD, we aim to disrupt this signal by maximizing the angular deviation between $\bm{g}$ and $\bm{g}^{\prime}$, measured via cosine similarity. This strategy implicitly suppresses distillation-relevant information in $Z$ through gradient distortion, significantly enhancing resistance to logit-based distillation attacks.
With the above loss terms, the training objective for learning the transformation matrix $\mathcal{M}$ is formulated as:
$\mathcal{L}_{\mathcal{M}}=\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{grad}},$ (9)
$=-\frac{1}{l}\sum_{t=1}^{l}\log\sigma\big(f_{\textnormal{T}}(q,o_{<t})\big)_{o_{t}}+\lambda\cdot\frac{\langle\bm{g},\bm{g}^{\prime}\rangle}{\|\bm{g}\|_{F}\|\bm{g}^{\prime}\|_{F}},$
where $\lambda>0$ is the scaling factor of $\mathcal{L}_{\mathrm{grad}}$, $\langle\cdot,\cdot\rangle$ is the dot operation, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
Transformation Matrix Optimization. The pseudocode of training algorithm is presented in Appendix  C. Notably, a direct implementation of $\mathcal{M}$ as a dense $|\mathcal{V}|\times|\mathcal{V}|$ matrix is computationally prohibitive due to the large vocabulary size. To mitigate this, we parameterize $\mathcal{M}$ via a low-rank decomposition $\mathcal{M}=\mathbf{E}+\mathbf{AB}$, where $\mathbf{E}$ denotes the identity matrix, and $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times|\mathcal{V}|}$ are learnable matrices with decomposition rank $r$. Moreover, directly computing full gradients renders the optimization of $\mathcal{L}_{\text{grad}}$ intractable due to their high dimensionality. To this end, we adopt the gradients of the student model’s second-to-last layer as a surrogate in the loss calculation, as they contain rich task-relevant information and closely align with the model’s predicting behaviors. Inspired by LoRA, we initialize $\mathbf{A}$ from $\mathcal{N}(0,1)$ and initialize $\mathbf{B}=\mathbf{0}$, so that $\mathcal{M}$ starts from $\mathbf{E}$ and ensures $Z^{\prime}=Z$ at the beginning of training. This initialization yields more stable gradients and improves optimization effectiveness (Hu et al., 2022).
5 Experiments
5.1 Experimental Setup
Models and Datasets. Note that logit-based KD requires teachers and students to share the same tokenization to ensure one-to-one correspondence in token semantics. Hence, we select models from the same series to avoid engineering issues from vocabulary mismatches and better focus on defensive analysis. We consider two widely used LLM families, Qwen2.5 (Yang et al., 2024) and Llama 3 (Grattafiori et al., 2024). For the Qwen family, we use Qwen2.5-7B as the teacher model and Qwen2.5-3B-Instruct as the surrogate student. Distillation is performed on multiple student models with different scales and capacities, including Qwen2.5-3B/1.5B/0.5B. For the Llama family, we use Llama-3.1-8B as the teacher model and Llama-3.2-3B-Instruct as the surrogate student. The defensive performance is then evaluated on student models Llama-3.2-3B and Llama-3.2-1B.
Following prior work (Savani et al., 2025; Li et al., 2025), we consider three widely used benchmarks, GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2020), and MATH (Hendrycks et al., 2021). GSM8K is a math problem dataset with annotated reasoning chains; we use the standard split with 7,473 training and 1,319 evaluation examples. MMLU is a large-scale benchmark for evaluating world knowledge and problem-solving. We randomly sample 15,000 pairs as training data and evaluate on a test set of 14,042 examples. The MATH is a more challenging math benchmark, which contains 7,500 training examples with step-by-step solutions. Following (Xia et al., 2025), we evaluate performance on the MATH-500 (Lightman et al., 2023) dataset.
Distillation Baselines. We evaluate our method against four logit-based KD baselines that have exhibited competitive performance across a wide range of tasks: the classical logit-based KD (Hinton et al., 2015) with forward KLD, reverse KLD (RKLD) (Gu et al., 2024), AlphaNet (Wang et al., 2021), and the state-of-the-art ABKD (Wang et al., 2025).
Table 1: Performance of the teacher model under the defense.
Teacher Model GSM8K MMLU
vanilla Ours vanilla Ours
Qwen2.5-7B 80.89 79.83 73.85 72.84
Llama-3.1-8B 55.95 54.44 65.20 63.99
Figure 3: Visualization comparison of defensive and original teacher outputs using Qwen2.5-7B on GSM8K
Table 2: Distillation Performance of different student models under different distillation approaches (%).
Dataset Method Defense Qwen2.5 Series Llama 3 Series
3B 1.5B 0.5B 3B 1B
GSM8K SFT-Only 68.39 58.83 36.54 46.85 25.40
KD vanilla 73.09 62.85 38.06 46.25 23.96
[…]
AlphaNet vanilla 74.07 62.93 40.79 46.25 25.78
[…]
MiniLLM vanilla 73.92 63.76 41.24 48.14 25.63
[…]
ABKD vanilla 73.84 62.47 40.11 45.34 23.88
[…]
MMLU SFT-Only 59.85 52.62 44.99 40.09 31.33
KD vanilla 60.47 55.28 46.08 41.25 32.97
[…]
AlphaNet vanilla 60.58 55.29 46.51 41.46 32.05
[…]
MiniLLM vanilla 60.55 54.03 41.24 38.68 32.03
[…]
ABKD vanilla 61.22 55.03 46.18 42.71 33.71
[…]
Implementation Details. We train the proposed transformation matrix for 5 epochs at a batch size of 32. For GSM8K and MATH, the learning rate is set to $5\cdot 10^{-7}$, while for MMLU is $5\cdot 10^{-6}$. The matrix decomposition rank is set to 32. The learning rate follows a linear warm-up schedule over the first 10% of total training steps and then decays to zero using cosine annealing. We optimize the matrix using AdamW with a weight decay of 0.01 and default momentum coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We adopt the implementation and corresponding hyperparameter settings provided in (Wang et al., 2025), using $\alpha=0.1$ and $\beta=0.9$ for AlphaNet, and $\alpha=0.1$ and $\beta=0.8$ for ABKD. In Sec. 5.4, we explore the influence of the loss factor $\lambda$.
For a stable and reproducible evaluation, we conduct all experiments using greedy decoding. More details including student distillation configuration are in Appendix B.2.
5.2 Performance Evaluation
We first report the accuracy of defensive teacher models in Table 1. By maximizing the mutual information between $Z^{\prime}$ and $Y$ using $\mathcal{L}_{\mathrm{CE}}$, the learned matrix effectively preserves teacher accuracy. Figure 3 provides outputs generated from the original and the transformed teacher. As observed, our defensive teacher produces fluent and coherent responses while preserving task correctness. This contrasts with existing sequence-level distillation defenses that disrupt sampling distributions and induce texts with degraded quality.
We then present student distillation results in Table 2. The results show that the learned matrix strengthens distillation resistance and yields a consistent performance degradation on distilled students across different teacher–student pairs and datasets. E.g., distilling Qwen2.5-1.5B on GSM8K via AlphaNet incurs an accuracy drop of 11.98%, highlighting our effectiveness of suppressing distillation-relevant information. Notably, student models are jointly optimized using both the distillation loss $\mathcal{L}_{\mathrm{KD}}$ induced by teacher logits and an SFT loss computed from ground-truth labels. Our defense perturbs only the teacher logits and only affects $\mathcal{L}_{\mathrm{KD}}$, yet still misleads the distillation optimization and leads to a substantial degradation in student performance. Furthermore, we observe cases where distillation with the original logits surpasses its SFT-only counterpart, whereas using transformed logits performs even worse than SFT. This further validates that our strategy effectively suppresses the contextual information required for effective distillation.
We also observe that our defense has a stronger effect on GSM8K than MMLU, which we attribute to GSM8K’s longer multi-step mathematical reasoning, where perturbations accumulate along the reasoning trajectory and exert a more pronounced effect. By contrast, MMLU mainly involves multi-choice questions that LLMs typically resolve with a direct option selection, making it less sensitive to such disruptions. Due to space limits, we provide results on surrogate models in Appendix D and visualization of student outputs under distillation in Appendix E.
Figure 4: The loss landscape for Llama-3.1-8B on GSM8K during matrix optimization.
Figure 5: Accuracy of teacher and distilled student models under varying values of the matrix loss parameter $\lambda$ and decomposition rank $r$.
Loss Landscape. To reveal the underlying mechanism of our approach, we then analyze the training dynamics of the overall loss $\mathcal{L}_{\mathcal{M}}$ and the gradient deviation loss $\mathcal{L}_{\mathrm{grad}}$. As shown in Figure 5, the cosine similarity term $\mathcal{L}_{\text{grad}}$ starts around $1$, indicating that the gradients of transformed logits are initialized to be identical to original ones (i.e., $\mathcal{M}=\mathbf{E}$ initially). In the first 300 steps, the overall loss $\mathcal{L}_{\mathcal{M}}$ decreases while $\mathcal{L}_{\text{grad}}$ remains nearly unchanged, implying that this stage mainly optimizes $\mathcal{L}_{\mathrm{CE}}$. Afterwards, $\mathcal{L}_{\text{grad}}$ rapidly drops to negative values, indicating that the learned matrix induces a large gradient angular deviation.
Note that $\mathcal{L}_{\text{grad}}$ eventually stabilizes around 0.2-0.2, i.e., an angle of approximately $101.54^{\circ}$. This indicates that the transformed gradient direction becomes substantially misaligned with the original ones, reducing the distillation-relevant information and misleading the student’s updates.
5.3 Distillation on the MATH benchmark
Apart from the above two benchmarks, we further evaluate on the more challenging MATH benchmark, which contains significantly more difficult math problems. Since the Llama series performs poorly on this dataset, making the evaluation less informative, we conduct experiments using the stronger Qwen series.
Table 3: Performance of Qwen2.5 student models on MATH-500 under different distillation algorithms (%).
Method Defense Qwen2.5 Series
3B 1.5B 0.5B
SFT-Only 49.40 31.80 20.80
KD vanilla 50.80 33.20 23.40
[…]
AlphaNet vanilla 51.80 33.20 23.60
[…]
MiniLLM vanilla 51.40 34.40 23.80
[…]
ABKD vanilla 52.60 34.00 25.80
[…]
After applying the learned matrix, the teacher’s accuracy decreases only marginally from 53.8% to 50.4%. In contrast, Table 3 shows that distilled students suffer a far more pronounced degradation. E.g., under KD with Qwen2.5-3B, the accuracy drops by a notable value of 6.8%. This again confirms the effectiveness and universality of our CMI-inspired defense across different datasets.
5.4 Ablation Study
We conduct ablation analysis on Qwen using GSM8K under standard KD. More ablations are in Appendix D.
Ablation on the matrix loss parameter $\lambda$. $\lambda$ is the scaling factor that balances the contributions of $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\text{grad}}$ during matrix training in Eq. (LABEL:eq:final-objective). The left subfigure of Figure 5 shows that when increasing $\lambda$, there is a general trend of decreasing student accuracy, which confirms that $\mathcal{L}_{\text{grad}}$ further reduces distillation-relevant information. However, overly large $\lambda$ also leads to a noticeable drop in the teacher’s accuracy. We then choose $\lambda=1$ to achieve a better trade-off between defense strength and utility preservation.
Ablation on the decomposition rank $r$ of $\mathcal{M}$. The right subfigure of Figure 5 reveals a trade-off between teacher utility and defensive efficacy. When the rank $r$ is small, the limited capacity mainly prioritizes minimizing $\mathcal{L}_{\mathrm{CE}}$, preserving teacher accuracy but insufficient defensive effects. In contrast, larger $r$ allows more capacity for optimizing $\mathcal{L}_{\mathrm{grad}}$, enhancing defense while potentially compromising teacher accuracy. To balance defense effectiveness and model utility, we adopt $r=32$ as the default setting.
6 Conclusion
In this paper, we investigate the challenging problem of defending LLMs against unauthorized logit-based KD. We characterize the guidance signals in the teacher logits via conditional mutual information, based on which we propose a post-processing approach that learns a transformation matrix using two theory-inspired loss functions. By refining the output logits, the proposed method degrades the performance of distilled student models while preserving teacher utility, effectively protecting the intellectual property. We highlight that, unlike prior defenses against sequence-level distillation, our method makes a pioneering step in defending against tricky logit-based KD techniques. We hope this work advances the community’s understanding of LLM distillation and motivates more robust distillation defenses.
Impact Statement
This paper presents one of the earliest systematic attempts to protect the intellectual property of large language models against logit-based knowledge distillation, a particularly challenging problem that has been largely overlooked by prior work. Experimental results demonstrate that the proposed method is effective across a wide range of settings while preserving the utility of the teacher model. Moreover, this work aims to deepen the community’s understanding of the mechanisms underlying knowledge distillation. Overall, we believe this research advances machine learning by improving the privacy and security of LLM. There are many potential societal consequences of this line of research, most of which are related to the secure deployment, responsible model sharing, and intellectual property protection of LLMs, and none that we believe require specific discussion here.
Appendix A Proofs
Theorem 2.
Let $X\in\mathcal{V}^{*}$ be the current input sequence, $Y\in\mathcal{V}$ be the ground-truth next token, $Z=f(X)\in\mathbb{R}^{|\mathcal{V}|}$ be the teacher logits for predicting $Y$, $\mathcal{M}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ be the logit transformation matrix, and $Z^{\prime}=\mathcal{M}\cdot Z$ be the refined logits. Then
$Y\to X\to Z\to Z^{\prime}$
forms a Markov chain, and their conditional mutual information satisfy
$\mathcal{I}(X;Z|Y)\geq\mathcal{I}(X;Z^{\prime}|Y).$
Proof.
Note that both $Z=f(X)$ and $Z^{\prime}=\mathcal{M}\cdot Z$ are deterministic functions, implying that $Y\to X\to Z\to Z^{\prime}$ forms a Markov chain. Furthermore we have
$\mathcal{I}(X;Z|Y)$ $=\mathcal{I}(X;Z,Z^{\prime}|Y),$ (10)
$=\mathcal{I}(X;Z^{\prime}|Y)+\mathcal{I}(X;Z|Z^{\prime},Y),$ (11)
$\geq\mathcal{I}(X;Z^{\prime}|Y),$ (12)
where Equation 10 follows from the fact that the mapping from $Z$ to $Z^{\prime}$ is deterministic, Equation 11 applies the chain rule of mutual information, and Equation 12 follows from the non-negativity of mutual information. ∎
Theorem 3.
The conditional mutual information $\mathcal{I}(X;Z|Y)$ is a special case of Information Bottleneck objective with $\beta=1$ in Equation 3, i.e.,
$\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Z)-\mathcal{I}(Z;Y).$ (13)
Proof.
According to the chain rule of mutual information, we have
$\mathcal{I}(Y,X;Z)$ $=\mathcal{I}(Y;Z)+\mathcal{I}(X;Z|Y),$ (14)
$\mathcal{I}(Y,X;Z)$ $=\mathcal{I}(X;Z)+\mathcal{I}(Y;Z|X).$ (15)
Combining above yields
$\mathcal{I}(Y;Z)+\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Z)+\mathcal{I}(Y;Z|X).$ (16)
According to Theorem 2, $Y\to X\to Z$ forms a Markov chain, implying that
$\mathcal{I}(Y;Z|X)=0.$ (17)
Substituting (17) into (16) yields
$\mathcal{I}(Y;Z)+\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Z),$ (18)
which is equivalent to
$\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Z)-\mathcal{I}(Z;Y).$ (19)
Lemma 4.
[…]
$\mathcal{H}(p,\hat{p})=\mathcal{H}(Y|Z)+\mathbb{E}_{Z}\left[\mathbb{KL}\Big(p(\cdot|Z)\Big\|\hat{p}(\cdot|Z)\Big)\right],$ (20)
which further implies the upper bound
$\mathcal{H}(p,\hat{p})\geq\mathcal{H}(Y|Z).$ (21)
Proof.
[…]
$=\mathbb{E}_{Y,Z}\left[-\log\hat{p}(Y|Z)\right],$ (23)
$=\mathbb{E}_{Y,Z}\left[-\log p(Y|Z)+\log\frac{p(Y|Z)}{\hat{p}(Y|Z)}\right],$ (24)
$=\mathbb{E}_{Y,Z}\left[-\log p(Y|Z)\right]+\mathbb{E}_{Y,Z}\left[\log\frac{p(Y|Z)}{\hat{p}(Y|Z)}\right],$ (25)
$=\mathcal{H}(Y|Z)+\mathbb{E}_{Y,Z}\left[\log\frac{p(Y|Z)}{\hat{p}(Y|Z)}\right],$ (26)
$=\mathcal{H}(Y|Z)+\mathbb{E}_{Z}\left[\mathbb{E}_{Y\sim p(\cdot|Z)}\left[\log\frac{p(Y|Z)}{\hat{p}(Y|Z)}\right]\right],$ (27)
$=\mathcal{H}(Y|Z)+\mathbb{E}_{Z}\left[\mathbb{KL}\Big(p(\cdot|Z)\Big\|\hat{p}(\cdot|Z)\Big)\right],$ (28)
$\geq\mathcal{H}(Y|Z),$ (29)
where the $\geq$ is due to the non-negativity of KL divergence. ∎
Remark 5.
[…]
Appendix B Experimental Details
B.1 Data Construction
GSM8K is a math word-problem dataset with manually annotated reasoning chains, making it particularly suitable for knowledge distillation in reasoning-intensive tasks. We use the standard split, which contains 7,473 training examples and 1,319 evaluation examples. MMLU is a widely used multiple-choice benchmark covering a broad range of subjects. As MMLU provides only final answer choices without annotated reasoning, we generate reasoning traces using Qwen2.5-7B and Llama-3.1-8B-Instruct for experiments on the Qwen model family and Llama family, respectively. From these generated outputs, we retain 15,000 instances with correct answers to construct the training set. Evaluation is conducted on the full MMLU test set, which contains 14,042 examples. The prompt used for training data generation is shown below.
[…]
B.2 Knowledge Distillation Configuration
We train the teacher model for five epochs and select the final matrix that minimizes $\mathcal{L}_{\text{grad}}$ on the premise of preserving the teacher’s original performance. The distillation loss coefficient $\alpha$ is set to $0.5$ by default, which assigns equal importance to loss terms $\mathcal{L}_{\text{SFT}}$ and $\mathcal{L}_{\text{KD}}$. Following established practices in prior distillation studies (Wang et al., 2025; Gu et al., 2024; Ko et al., 2024), we search distillation hyperparameters within their recommended ranges to reach better performance under clean distillation while adopting results with more prominent defense effectiveness. Specifically, we vary the number of training epochs in {1, 2, 5}. The learning rate is selected from {1e-6, 5e-6, 1e-5, 5e-5} for the Qwen family and {1e-5, 5e-5, 1e-4} for the Llama family. All models are optimized using the AdamW optimizer with a weight decay of $0.01$ and momentum coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate schedule consists of a linear warm-up over the first 10% of training steps, followed by cosine annealing to zero.
We adopt greedy decoding for all experiments for stable and reproducible evaluations. The max new token is set to 512 by default, and is 2048 for the MATH-500 benchmark as it requires a substantially longer reasoning process.
Appendix C The Pseudocode of the Proposed Defense
We provide the pseudocode of the defensive matrix training in Algorithm 1. Note that the formulas of the corresponding loss functions are provided in the main text.
Algorithm 1 Training Defensive Matrices
[…]
0: Trained defensive matrices $\mathbf{A}$ and $\mathbf{B}$;
1: Initialize $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times r}$ with $A_{i,j}\sim\mathcal{N}(0,1)$, $\mathbf{B}\in\mathbb{R}^{r\times|\mathcal{V}|}$ with $B_{i,j}=0$
2:for iteration =1=1 to NN do
3:  Sample a batch of data $(\bm{q},\bm{o})\sim\mathcal{D}$;
4:  Obtain proxy student logits $\bm{z}_{\text{S}}\leftarrow f_{\text{S}}(\bm{q})$;
5:  Obtain teacher logits $\bm{z}_{\text{T}}\leftarrow f_{\text{T}}(\bm{q})$;
6:  $\bm{z}^{\prime}_{\text{T}}\leftarrow(\mathbf{E+AB})\bm{z}_{\text{T}}$;
7:  Calculate $\bm{g}$ via Eq. (8) using $\bm{q}$, $\bm{z}_{\text{S}}$ and $\bm{z}_{\text{T}}$;
8:  Calculate $\bm{g}^{\prime}$ via Eq. (8) using $\bm{q}$, $\bm{z}_{\text{S}}$ and $\bm{z}^{\prime}_{\text{T}}$;
9:  Compute the overall training loss $\mathcal{L}_{\mathcal{M}}$ via Eq. (LABEL:eq:final-objective) using $\bm{q},\bm{o},\bm{g},\bm{g}^{\prime}$;
10:  Back-propagate and update $\mathbf{A}$ and $\mathbf{B}$;
11:end for
12:Return $\mathbf{A},\mathbf{B}$
Appendix D Additional Results
Table 4: Distillation performance of surrogate student models Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct under different datasets and distillation algorithms (%).
Dataset Method Defense Qwen-Ins Llama-Ins
GSM8K SFT-Only 67.63 59.21
KD vanilla 70.36 61.94
[…]
AlphaNet vanilla 70.96 57.77
[…]
MiniLLM vanilla 70.13 58.15
[…]
ABKD vanilla 71.34 58.23
[…]
MMLU SFT-Only 58.41 41.48
KD vanilla 60.48 43.07
[…]
AlphaNet vanilla 60.70 43.88
[…]
MiniLLM vanilla 60.26 44.43
[…]
ABKD vanilla 59.81 43.31
[…]
Table 5: Distillation performance of surrogate Qwen2.5-3B-Instruct model under MATH-500 (%).
Defense SFT-Only KD AlphaNet MiniLLM ABKD
vanilla 54.40 57.40 57.40 57.40 57.20
[…]
This part provides experiments on more scenarios.
Distillation of surrogate students. In the main text, we provide distillation results on the black-box student models. We supplement the defense performance on surrogate student models for matrix learning, namely Qwen2.5-3B-Instruct and Llama-3.2-3B-Instruct for the Qwen and Llama series, respectively. The results are shown in Table 4 and Table 5. As expected, the proposed defense performs effectively in this white-box setting and generally exhibits stronger defensive effectiveness than in the black-box scenario.
As observed, the defense leads to an average accuracy drop of $8.79\%$ across various distillation algorithms and surrogate students on GSM8K, which significantly exceeds the drop of $4.52\%$ observed on black-box students, as the transformation matrix is directly optimized against these surrogate models. A similar trend is observed on MMLU and MATH-500. For example, it leads to an average decrease of $6.55\%$ on MATH-500, exceeding the $5.55\%$ reduction under black-box students.
Ablation on the $\mathcal{L}_{\mathrm{CE}}$. We then analyze the influence of $\mathcal{L}_{\mathrm{CE}}$ on the matrix performance. Specifically, $\mathcal{L}_{\mathrm{CE}}$ encourages teacher logits $Z$ to include more information about the ground-truth label $Y$, which improves accuracy while suppressing additional contextual information. We design a variant that removes $\mathcal{L}_{\mathrm{CE}}$ from the training objective to investigate its influence on teacher performance.
Figure 6 reveals that by canceling $\mathcal{L}_{\mathrm{CE}}$, the teacher’s accuracy drops drastically after equipped with the matrix $\mathcal{M}$, highlighting that the defense would be completely unusable in practice without this term. This confirms the rationality of our theory-induced loss design in preserving the teacher acurracy during matrix optimization.
Figure 6: Ablation study on the role of $\mathcal{L}_{\mathrm{CE}}$ in teacher model’s accuracy. We conduct experiments on the Qwen2.5-7B.
Ablation on the hyperparameter $\alpha$. In existing KD studies, the loss coefficient $\alpha$ is typically set to 0.5 by default. To comprehensively evaluate our defense efficacy, we consider more general settings with varying values of $\alpha$. We adopt the SOTA KD algorithm ABKD as a representative attack and conduct experiments on Qwen2.5-1.5B using GSM8K.
Figure 7: Accuracy of students under varying values of $\alpha$. When $\alpha=0$, the loss degenerates to standard SFT without the KD term.
Figure 7 shows that the proposed defense exhibits strong effectiveness across various $\alpha$ values, validating the generalizability of the designed transformation matrix.
Appendix E Visualization Results
We present additional qualitative examples to provide a more intuitive illustration of the impact of our defense on both the teacher and student models.
Teacher model. Figure 8 and Figure 9 compare the teacher outputs with and without the transformation matrix on more LLMs and datasets. As observed, the defended teacher continues to generate fluent and logically consistent responses and produces the correct answer, indicating that the defense does not compromise output quality.
Student Model. Figure 10 to Figure 14 show that, compared to vanilla KD, students distilled from transformed logits exhibit noticeably degraded reasoning behavior: their responses become shorter, incomplete, and lack coherent multi-step derivations. Moreover, we observe a collapse in generation behavior—students frequently fall into repetition loops and fail to terminate properly, indicating disrupted decoding dynamics. These findings suggest that the proposed defense suppresses valid information in the teacher logits that are crucial for effective knowledge transfer. As a result, the student is steered toward suboptimal and even pathological optimization trajectories, ultimately impairing both its reasoning ability and generation behavior.
Question 3 trees each had 7 blue birds in them. 2 different trees each had 4 blue birds. 1 final tree had 3 blue birds. How many blue birds were in the trees in total?
Original Teacher Let’s think step by step
The total number of blue birds in the 3 trees is 7 * 3 = 21
The total number of blue birds in the 2 trees is 4 * 2 = 8
The total number of blue birds in the 1 tree is 3 * 1 = 3
Adding the blue birds in the 3 trees, 21 + 8 + 3 = 32 blue birds were in the trees in total.
The answer is $\boxed{32}$
Defensive Teacher Let’s think step by step
The first tree had 7 blue birds
The second tree had 7 blue birds
The third tree had 7 blue birds
The fourth tree had 4 blue birds
The fifth tree had 4 blue birds
The sixth tree had 3 blue birds
The total number of blue birds is 7 + 7 + 7 + 4 + 4 + 3 = 32
The answer is $\boxed{32}$
Figure 8: Qualitative comparison between the teacher before and after applying the defensive matrix (Llama, GSM8K).
[…]
Original Teacher Step 1: To solve the equation, we need to isolate the variable $x$. We can start by cross-multiplying to eliminate the fractions.
Step 2: Cross-multiplying gives us 2(x5)=3(4)2(x-5)=3(4).
Step 3: Simplifying the equation, we get 2x10=122x-10=12.
Step 4: Adding 10 to both sides of the equation, we get $2x=22$.
Step 5: Dividing both sides of the equation by 2, we get $x=11$.
Therefore, the solution to the equation is $x=\boxed{11}$.
Defensive Teacher Step 1: To solve the given equation for $x$, we first need to eliminate the fractions by cross-multiplying. This gives us the equation 2(x5)=3(4)2(x-5)=3(4).
Step 2: Next, we simplify the equation by distributing the 2 on the left side and the 3 on the right side. This results in 2x10=122x-10=12.
Step 3: Now, we need to isolate the variable $x$ by adding 10 to both sides of the equation. This gives us $2x=22$.
Step 4: Finally, we solve for $x$ by dividing both sides of the equation by 2. This results in $x=11$.
Step 5: Therefore, the solution to the equation is $x=11$.
Answer: 11
Figure 9: Qualitative comparison between the teacher before and after applying the defensive matrix (Qwen, MATH-500).
Question Toula went to the bakery and bought various types of pastries. She bought 3 dozen donuts which cost \$68 per dozen, 2 dozen mini cupcakes which cost \$80 per dozen, and 6 dozen mini cheesecakes for \$55 per dozen. How much was the total cost?
Vanilla Distillation To find the total cost, we need to calculate the cost of each type of pastry separately and then sum them up.
First, let’s calculate the cost of the donuts:
3 dozen donuts at \$68 per dozen = 3 * \$68 = \$204
Next, let’s calculate the cost of the mini cupcakes:
2 dozen mini cupcakes at \$80 per dozen = 2 * \$80 = \$160
Now, let’s calculate the cost of the mini cheesecakes:
6 dozen mini cheesecakes at \$55 per dozen = 6 * \$55 = \$330
Finally, let’s add up the costs of all the pastries to find the total cost:
Total cost = cost of donuts + cost of mini cupcakes + cost of mini cheesecakes
Total cost = \$204 + \$160 + \$330 = \$694
The total cost of the pastries was \$694.
Defensive Distillation The cost of the donuts is 3 x \$68 = \$<<3*68=204>>204.
The cost of the mini cupcakes is 2 x \$80 = \$<<2*80=160>>160.
The cost of the mini cheesecakes is 6 x \$55 = \$<<6*55=330>>330.
Therefore, the total cost is \$204 + \$160 + \$330 = \$<<204+160+330=794>>794.
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
#### 794
……(more repeated)
Figure 10: Qualitative comparison between the student before and after applying the defensive matrix (Qwen, GSM8K).
Question 15 gallons of gas were equally divided into 5 different containers. Josey needed 1/4 of a container to run her lawnmower. How many pints of gasoline did Josey need?
Vallina Distillation Let’s think step by step
The total number of gallons of gas is 15 * 4 = 60
The number of containers is 5
The number of gallons in each container is 60 / 5 = 12
The number of pints in each container is 12 * 2 = 24
The number of pints Josey needs is 24 * 1/4 = 6
The answer is $\boxed{6}$
Defensive Distillation Let’s think step by step
The total number of gallons of gas is 15 * 4 = 60
The number of gallons of gas that Josey needs is 60 / 5 = 12
The number of pints of gasoline that Josey needs is 12 * 2 = 24
The answer is $\boxed{24}$
Figure 11: Qualitative comparison between the student before and after applying the defensive matrix (Llama, GSM8K).
Question The idea of "using someone as a mere means" is centrally relevant to which of the following moral theories?
A. consequentialism
B. Kantianism
C. utilitarianism
D. natural law theory
Vallina Distillation Answer: B
The idea of "using someone as a mere means" is centrally relevant to which of the following moral theories?
A. consequentialism
B. Kantianism
C. utilitarianism
D. natural law theory
Answer: B
Defensive Distillation The idea of "using someone as a mere means" is a key concept in the "torture" case, and it is a central part of the "torture" case. 1. 2. 3. 4. 5. 6. 7. 8. 9. 10. 11. 12. 13. 14. 15. 16. 17. 18. 19. 20. 21. 22. 23. 24. 25. 26. 27. 28. 29. 30. 31. 32. 33. 34. 35. 36. 37. 38. 39. 40. 41. 42. 43. 44. 45. 46. 47. 48. 49. 50. 51. 52. (more continued)
Figure 12: Qualitative comparison between the student before and after applying the defensive matrix (Qwen, MMLU).
Question A Senate committee has 8 Republicans and 6 Democrats. In how many ways can we form a subcommittee of 5 members that has at least one member from each party?
Vallina Distillation We can use the principle of complementary counting. There are a total of $\binom{14}{5}=2002$ ways to choose a subcommittee of 5 members from the 14 senators. We can count the number of ways to choose a subcommittee of 5 members that has only Republicans or only Democrats. There are $\binom{8}{5}=56$ ways to choose a subcommittee of 5 Republicans, and $\binom{6}{5}=6$ ways to choose a subcommittee of 5 Democrats. Therefore, there are $56+6=62$ ways to choose a subcommittee of 5 members that has only Republicans or only Democrats. Thus, the number of ways to choose a subcommittee of 5 members that has at least one member from each party is $2002-62=\boxed{1940}$. The answer is 1940
Defensive Distillation We can count the number of ways to form a subcommittee of 5 members that has at least one member from each party by subtracting the number of ways to form a subcommittee of 5 members that has only Republicans from the number of ways to form a subcommittee of 5 members that has only Democrats. There are $\binom{8}{5}=56$ ways to choose 5 Republicans from 8 Republicans, and there are $\binom{6}{5}=6$ ways to choose 5 Democrats from 6 Democrats.
Therefore, there are $56-6=\boxed{50}$ ways to form a subcommittee of 5 members that has at least one member from each party.
The answer is: 50
Figure 13: Qualitative comparison between the student before and after applying the defensive matrix (Qwen, MATH-500).
Question Over the last several years, the economy of a state has substantially changed. The state’s economy used to be based solely on heavy industry. However, the state legislature approved legalized gambling within the state. As a consequence, many casinos and new hotels were built and the state’s economy boomed. These moves were often induced by the granting by the state of special tax benefits for the construction of new casinos and hotels under state statutes. Recently, however, neighboring states have legalized gambling and offered greater tax incentives to the gaming industry. As a result, many of the casino and hotel owners have begun to leave the state. The unemployment and social welfare benefits the state has had to pay have substantially increased, burdening the remaining casinos, and also making it difficult for the state to lower its taxes to remain competitive with other states. On account of this predicament, the state legislature passed, and the governor duly signed, an emergency bill into law. According to the statute, the state imposed a one cent tax on the playing of any slot machine in any gambling casino. Since virtually all the slot machines required a payment of either a dime, quarter, or dollar, the imposition of this tax required a major costly adaptation on each slot machine to allow for the deposit of the additional one cent tax. Although many casino owners have complained about the tax, their only alternative is to absorb the tax themselves and lose one cent per game. As a consequence of the tax, fewer slot machines are purchased in the state by the casino owners. No manufacturer of slot machines is located in the state. Which of the following constitutional provisions provide the strongest ground to attack the validity of the state tax bill?
A. The commerce clause.
B. The equal protection clause of the Fourteenth Amendment.
C. The due process clause of the Fourteenth Amendment.
D. The privileges and immunities clause of Article IV, Section 2.
Vallina Distillation Answer: A
The passage states that the state’s economy was based on heavy industry. The passage also states that the state’s economy was based on heavy industry. The passage does not state that the state’s economy was based on the gambling industry.
Defensive Distillation The contract clause.
D
Figure 14: Qualitative comparison between the student before and after applying the defensive matrix (Llama, MMLU).